OpenStudy (anonymous):

Simplify the following expression. 2a+3-9b-a-7

OpenStudy (anonymous):

a-9b-4

OpenStudy (anonymous):

just combining like terms, 2a- a = a 3-7= -4 -9b=-9b

OpenStudy (anonymous):

you cant mix a's with b's

OpenStudy (anonymous):

and you cant add integers( ie 3 and - 7) with the a's or b's

OpenStudy (anonymous):

so, to sum up, start from left to right, the first term you run into is 2a, then you pass by +3 -9b and run into -a

OpenStudy (anonymous):

so what is 2a - a?